# bibliography-2.bib

### v5

@@ Introduction
In his groundbreaking paper [Na52], Nash proved that every compact smooth manifold $M$ of dimension $d$ is diffeomorphic to a real analytic submanifold $M^{\prime}\subset{\mathbb{R}}^{2d+1}$ which is actually the union of some nonsingular connected components of a real algebraic subset of ${\mathbb{R}}^{2d+1}$. Hence, Nash conjectured that $M^{\prime}$ could be chosen to be a whole nonsingular real algebraic subset of ${\mathbb{R}}^{2d+1}$, a so-called algebraic model of $M$. In 1973, Tognoli [Tog73] proved this conjecture to be true, thus the Nash-Tognoli theorem asserts that: Every compact manifold $M$ of dimension $d$ is diffeomorphic to a nonsingular real algebraic subset of ${\mathbb{R}}^{2d+1}$. With respect to Nash’s result two main original ideas appear in Tognoli’s work: relative algebraic approximation of smooth functions by polynomial (and regular) functions and cobordism theory, in particular the algebraic representatives of cobordism classes produced by Milnor in [Mil65].
After Tognoli’s solution of Nash’s conjecture a systematic study of real algebraic sets and of the algebraicity problem started. There is a wide literature devoted to improvements and extensions of Nash-Tognoli theorem. We refer the interested reader to the books [AK92a, §II], [BCR98, §14], [Man14, §6], the survey [Kol17, §2], the papers [AK81b, AK81a, CS92, BK89, Kuc11], the more recent ones [Ben1, GT17] and references therein.
@@
For compact smooth manifolds, above problem for $k={\mathbb{R}}_{\mathrm{alg}}$ has an affirmative answer. This follows combining Nash’s theorem and the algebraicity result [CS92, Corollary 3.9] for Nash manifolds over any real closed field by Coste and Shiota. For $k={\mathbb{Q}}$ a positive answer is given by Ghiloni and the author in [GSa, Theorem 3.9]. Previously, Ballico and Tognoli [BT92, Theorem 0.1] stated a result which is even stronger but, unfortunately, the proof of this theorem is not complete, so the result is not guaranteed to be true. In Remark 4.3 we outline why some arguments of the classical Nash-Tognoli theorem do not directly extend to obtain polynomial equations over ${\mathbb{Q}}$. In particular, we stress that several subtleties arise when looking for equations over ${\mathbb{Q}}$ so a different notion of regularity over $\mathbb{Q}$ (see Definition 1.4) with respect of the one proposed in [Tog78, Definition 3, p. 30] and used in [BT92] is required to fill the gaps in the proof. However, we recognize to Ballico and Tognoli the original idea of extending the algebraic approximation techniques to obtain nonsingular equations with coefficients over ${\mathbb{Q}}$ for the algebraic models of compact smooth manifolds.
@@
We outline that both the approaches by Coste-Shiota and Parusiński-Rond make use of the model completeness of the theory of real closed fields. That is the reason why their techniques can not be extended to rational numbers, indeed ${\mathbb{Q}}$ is not a real closed field and the theory of ordered fields is not model complete. Hence, the unique approach which seems to work at the moment in investigating the ${\mathbb{Q}}$-algebraicity problem is the one proposed by Ghiloni and the author in [GSa] and which is further developed in the present paper.
In order to investigate latter open problem, as already mentioned, classical notions in algebraic geometry has to be very carefully adapted when looking for descriptions over ${\mathbb{Q}}$. In the brand new paper [FG], Fernando and Ghiloni introduced and studied ${\mathbb{Q}}$-algebraic subsets of ${\mathbb{R}}^{n}$, that is, real algebraic sets globally described by polynomial equations with rational coefficients. Let $V\subset{\mathbb{R}}^{n}$ be a ${\mathbb{Q}}$-algebraic set. Denote by ¯$\overline{{\mathbb{Q}}}:={\mathbb{R}}_{\mathrm{alg}}[i]$ the algebraic closure of ${\mathbb{Q}}$. Denote by $V_{\mathbb{C}}\subset{\mathbb{C}}^{n}$ the complexification of $V$ and choose some equations over ${\mathbb{R}}_{\mathrm{alg}}$ defining $V_{\mathbb{C}}$. Let $E|{\mathbb{Q}}$ be a finite Galois subextension of $\overline{{\mathbb{Q}}}|{\mathbb{Q}}$ containing all the coefficients of the chosen equations defining $V_{\mathbb{C}}$. By means of the action of the Galois group $\textnormal{Gal}(E|{\mathbb{Q}})$ on $V_{\mathbb{C}}$, they proved fundamental properties of real ${\mathbb{Q}}$-algebraic sets and defined a precise notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity for a point of a real ${\mathbb{Q}}$-algebraic set (see Definition 1.4). Latter notion of regularity over ${\mathbb{Q}}$ for a point on a real ${\mathbb{Q}}$-algebraic set turned out to be crucial to extend the algebraic approximation techniques developed in [Na52], [Tog73] and [AK81a]. Indeed, in [GSa] Ghiloni and the author developed such ${\mathbb{Q}}$-algebraic approximation techniques and gave a complete solution of above ${\mathbb{Q}}$-algebraicity problem in the case of nonsingular real algebraic sets and real algebraic sets with isolated singularities. As already mentioned, in [GSa] Ghiloni and the author proved, as a special case, the following version over ${\mathbb{Q}}$ of Nash-Tognoli theorem: Every compact manifold $M$ of dimension $d$ is diffeomorphic to a ${\mathbb{Q}}$-nonsingular real ${\mathbb{Q}}$-algebraic subset of ${\mathbb{R}}^{2d+1}$.
@@
${\mathbb{Q}}$-algebraicity problem in low dimensions: Our Main Theorem below is expected to be deeply used to provide a complete positive answer to the ${\mathbb{Q}}$-algebraicity problem in low dimensions. The 1-dimensional case is just a consequence of [GSa, Theorem 5.1 & Corollary 5.2], in dimension $2$ and $3$ the problem is much more challenging. We refer to [BD81, AK92a] for more details on the topological characterization of polyhedra of dimension up to $3$ admitting an algebraic structure and to [CP00] for necessary conditions on local invariants in all dimensions.
${\mathbb{Q}}$-algebraic homology of manifolds: It was shown by Thom [Tho54] that not every ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of compact smooth manifolds can be represented by a closed smooth submanifold, but it can be represented by the fundamental class of a compact smooth manifold by a smooth mapping. As a consequence, Kucharz [Kuc05] showed that any ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of a compact Nash manifold can be represented by a semialgebraic arc-symmetric subset. On the other hand, there are many homology classes that can not be represented by algebraic subsets (even singular), see [BD84, AK85, BK89, Ben1]. It is interesting then to see whether those classes that can be represented by real algebraic subsets can be realized by ${\mathbb{Q}}$-algebraic subsets.
@@
Section 1 is devoted to review the fundamental results of ${\mathbb{Q}}$-algebraic geometry developed in [FG, GSa]. We recall the notions of ${\mathbb{Q}}$-algebraic set, the decomposition in ${\mathbb{Q}}$-irreducible components and the notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity of a point in a ${\mathbb{Q}}$-algebraic set $V\subset{\mathbb{R}}^{n}$ (see Definition 1.2). This notion turned out to be crucial in separating the irreducible components of ${\mathbb{Q}}$-algebraic sets (see Corollary 1.6) and to develop ${\mathbb{Q}}$-algebraic approximation techniques in [GSa, §3]. In the last part of this first section we propose fundamental examples of ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic sets (see Definition 1.4) that became crucial in Subsection 3.1.
@@
Section 2 is devoted to the study of ${\mathbb{Z}}/2{\mathbb{Z}}$-homology groups of real Grassmannians. It is well known that each real Grassmannian ${\mathbb{G}}_{m,n}$ of affine $m$-planes in ${\mathbb{R}}^{m+n}$ can be embedded as an algebraic subset of ${\mathbb{R}}^{(m+n)^{2}}$, see [BCR98, Theorem 3.4.4]. Let us identify ${\mathbb{G}}_{m,n}$ with the above algebraic subset of ${\mathbb{R}}^{(m+n)^{2}}$. It is also well known that incidence conditions with respect to a complete flag induce a finite cellular decomposition of ${\mathbb{G}}_{m,n}$ such that the Euclidean closure of each cell is an algebraic subset of ${\mathbb{G}}_{m,n}$, see [MS74]. The closures of these cells are called Schubert varieties associated to the complete flag and they generate the ${\mathbb{Z}}/2{\mathbb{Z}}$-homology groups of each ${\mathbb{G}}_{m,n}$. The main result of this section is that, if we choose the complete flag $0\subset{\mathbb{R}}\subset{\mathbb{R}}^{2}\subset\dots\subset{\mathbb{R}}^{m+n}$ with respect to the standard coordinates of ${\mathbb{R}}^{m+n}$, each Schubert variety $\sigma_{\lambda}$ defined by incidence conditions, prescribed by a partition $\lambda$, with respect to the above complete flag admits a ${\mathbb{Q}}$-desingularization (see Definition 2.3). Latter result ensures that each ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of each real Grassmannian ${\mathbb{G}}_{m,n}$ has a ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic representative. This is a key property for the construction of ${\mathbb{Q}}$-algebraic relative bordisms in Section 3. We point out that very general algorithms in resolution of singularities in characteristic 00 provide a resolution which is functorial and invariant under field extension, see as references [Wlo05, Section 5.7], [Kol07, Section 3.34.2] and [BM08, Theorem 1.1]. However, mentioned algorithms have a very high complexity, thus it is a very hard task in general to concretely apply them obtaining an explicit resolution of a given algebraic variety, see [FKP]. By contrast, the application of the very explicit Bott-Samelson desingularization procedure for real embedded Schubert varieties, see [Dem74, Zel83] and the books [FP98, Man01, Bri05], allows us to conbinatorically control the equations defining the center of each blow-up of the resolution. We emphasize that, to the best of the author’s knowledge, the ${\mathbb{R}}|{\mathbb{Q}}$-regularity of real embedded Bott-Samelson varieties is not a direct consequence of results already known in the literature, although Demazure [Dem74] develops the desingularization procedure in the very general setting of flag varieties over a field $k$. Indeed, for our purposes, the choice of an appropriate embedding in some real affine space of each Grassmannian manifold ${\mathbb{G}}_{m,n}$ and of the resulting desingularization of its embedded Schubert subvarieties plays a crucial role for the notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity, see Example 1.3 for a clear explanation. We refer the interested reader to Remark 2.8 for more details on how to deduce the existence of a ${\mathbb{Q}}$-desingularization of ${\mathbb{Q}}$-algebraic sets from general results in resolution of singularities and why Bott-Samelson procedure is preferable in the case of Schubert varieties.
@@
Section 3 is divided in two different subsections. The first one is devoted to adapt ‘over ${\mathbb{Q}}$’ the topological construction of what we call “relative bordisms” introduced by Akbulut and King in [AK81b]. We stress that the topological ideas come from Akbulut and King paper but to ensure the equations to have rational coefficients, both globally and locally, our choice of an appropriate embedding is crucial. The second subsection is a review on ${\mathbb{Q}}$-nice ${\mathbb{Q}}$-algebraic sets introduced by Ghiloni and the author in [GSa].
@@ 1. Review on (real) ${\mathbb{Q}}$-algebraic geometry
1.1. ${\mathbb{Q}}$-Algebraic sets, ${\mathbb{R}}|{\mathbb{Q}}$-regularity and ${\mathbb{Q}}$-regular functions
In this subsection we briefly recall some fundamental notions of real and complex ${\mathbb{Q}}$-algebraic geometry introduced in [FG, GSa].
@@
$\mathcal{Z}_{L}(F)$ :=$:=\{x\in L^{n}:f(x)=0,\,\forall f\in F\},$
${\mathcal{I}}_{K}(S)$ :=$:=\{f\in K[x]:f(x)=0,\,\forall x\in S\}.$
@@
The $K$-dimension $\dim_{K}(V)$ of $V$ is defined as the Krull dimension of the ring $K[x]/{\mathcal{I}}_{K}(V)$. As above, $\dim_{L}(V)$ coincides with the usual dimension $\dim(V)$ of $V$, viewed as an algebraic subset of $L^{n}$. A $K$-version of a classical result concerning irreducibility and dimension holds: if $V\subset L^{n}$ is $K$-irreducible and $W\subset L^{n}$ is a $K$-algebraic set such that $W\subsetneq V$, then $\dim_{K}(W)<\dim_{K}(V)$. Moreover, if $L$ is algebraically closed or real closed, then $\dim_{K}(V)=\dim(V)$. For further information on these topics we refer to [FG, §1].
Moreover, if $L$ is algebraically closed, an application of Hilbert’s Nullstellensatz and Galois Theory in [FG, Theorem 1.3.4] guarantees that ${\mathcal{I}}_{L}(V)={\mathcal{I}}_{K}(V)L[x]$, thus the ideal ${\mathcal{I}}_{K}(V)$ of $K[x]$ gives complete information about $V\subset L^{n}$. Latter equality holds as well when both $L$ and $K$ are real closed fields by the Tarski-Seidenberg principle, see [FG, Lemma 1.4.1]. However, this is false for a general field extension $L|K$. As an example, if $L|K={\mathbb{R}}|{\mathbb{Q}}$ and $V:=\{x-\sqrt[3]{2}=0\}=\{x^{3}-2=0\}\subset{\mathbb{R}}$, then
@@
We say that an algebraic set $S\subset L^{n}$ is defined over $K$ if ${\mathcal{I}}_{L}(S)={\mathcal{I}}_{K}(S)L[x]$. Clearly, if $S\subset L^{n}$ is defined over $K$, then $S\subset L^{n}$ is also $K$-algebraic. As said before, if $L$ is algebraically closed or both $L$ and $K$ are real closed, then the concepts of $K$-algebraic subset of $L^{n}$ and algebraic subset of $L^{n}$ defined over $K$ do coincide. As explained in [GSa, §2.2], in the real algebraic setting the situation is completely different and various notions of nonsingularity over ${\mathbb{Q}}$ arise. In what follows we recall the definition of ${\mathbb{R}}|{\mathbb{Q}}$-regular points of a ${\mathbb{Q}}$-algebraic set introduced and studied in [FG] and whose characterizations are further investigated in [GSa].
@@
and by ${\mathcal{I}}_{\mathbb{Q}}(V)$ the vanishing ideal of $V$ in ${\mathbb{Q}}[x]$, as above. The following notion of ${\mathbb{R}}|{\mathbb{Q}}$-regular point was introduced in [FG, Definition 4.1.1].
@@
In [FG, Theorem 4.4.1], Fernando and Ghiloni show that the set $\operatorname{Reg}^{{\mathbb{R}}|{\mathbb{Q}}}(V)$ is a non-empty Zariski open subset of $\operatorname{Reg}(V)$. Though, it may happen that $\operatorname{Reg}^{{\mathbb{R}}|{\mathbb{Q}}}(V)\subsetneq\operatorname{Reg}(V)$, see the example below.
@@
[…]
This leads to the following definition originally introduced in [GSa, Definition 1.2].

### v6

# A relative Nash-Tognoli theorem over $\mathbb{Q}$ and application to the $\mathbb{Q}$-algebraicity problem

@@ Introduction
In his groundbreaking paper [Nas52], Nash proved that every compact smooth manifold $M$ of dimension $d$ is diffeomorphic to a real analytic submanifold $M^{\prime}\subset{\mathbb{R}}^{2d+1}$ which is actually the union of some nonsingular connected components of a real algebraic subset of ${\mathbb{R}}^{2d+1}$. Hence, Nash conjectured that $M^{\prime}$ could be chosen to be a whole nonsingular real algebraic subset of ${\mathbb{R}}^{2d+1}$, a so-called algebraic model of $M$. In 1973, Tognoli [Tog73] proved this conjecture to be true, thus the Nash-Tognoli theorem asserts that: Every compact manifold $M$ of dimension $d$ is diffeomorphic to a nonsingular real algebraic subset of ${\mathbb{R}}^{2d+1}$. With respect to Nash’s result two main original ideas appear in Tognoli’s work: relative algebraic approximation of smooth functions by polynomial (and regular) functions and cobordism theory, in particular the algebraic representatives of cobordism classes produced by Milnor in [Mil65].
After Tognoli’s solution of Nash’s conjecture a systematic study of real algebraic sets and of the algebraicity problem started. There is a wide literature devoted to improvements and extensions of Nash-Tognoli theorem. We refer the interested reader to the books [AK92, §II], [BCR98, §14], [Man14, §6], the survey [Kol17, §2], the papers [AK81, AK81a, CS92, BK89, Kuc11], the more recent ones [Ben24, GT17] and references therein.
@@
For compact smooth manifolds, above problem for $k={\mathbb{R}}_{\mathrm{alg}}$ has an affirmative answer. This follows combining Nash’s theorem and the algebraicity result [CS92, Corollary 3.9] for Nash manifolds over any real closed field by Coste and Shiota. For $k={\mathbb{Q}}$ a positive answer is given by Ghiloni and the author in [GS25, Theorem 1.7]. Previously, Ballico and Tognoli [BT92, Theorem 0.1] stated a result which is even stronger but, unfortunately, the proof of this theorem is not complete, so the result is not guaranteed to be true. In Remark 4.3 we outline why some arguments of the classical Nash-Tognoli theorem do not directly extend to obtain polynomial equations over $\mathbb{Q}$}. In particular, we stress that several subtleties arise when looking for equations over $\mathbb{Q}$} so a different notion of regularity over $\mathbb{Q}$ (see Definition 1.4) with respect to the one proposed in [Tog78, Definition 3, p. 30] and used in [BT92] is required to fill the gaps in the proof. In particular, the theory of subfield-algebraic geometry recently developed by Fernando and Ghiloni [FG25] plays an essential role to clarify the basic concepts involved and to produce useful tools for our purposes. However, we recognize to Ballico and Tognoli the original idea of extending the algebraic approximation techniques to obtain nonsingular equations with coefficients over $\mathbb{Q}$} for the algebraic models of compact smooth manifolds.
@@
We outline that both the approaches by Coste-Shiota and Parusiński-Rond make use of the model completeness of the theory of real closed fields. That is the reason why their techniques can not be extended to rational numbers, indeed $\mathbb{Q}$} is not a real closed field and the theory of ordered fields is not model complete. Hence, the unique approach which seems to work at the moment in investigating the $\mathbb{Q}$}-algebraicity problem is the one proposed by Ghiloni and the author in [GS25] and which is further developed in the present paper.
In order to investigate latter open problem, as already mentioned, classical notions in algebraic geometry must be very carefully adapted when looking for descriptions over $\mathbb{Q}$}. In the brand new paper [FG25], Fernando and Ghiloni introduced and studied $\mathbb{Q}$}-algebraic subsets of ${\mathbb{R}}^{n}$, that is, real algebraic sets globally described by polynomial equations with rational coefficients. Let $V\subset{\mathbb{R}}^{n}$ be a $\mathbb{Q}$}-algebraic set. Denote by $\overline{{\mathbb{Q}}}:={\mathbb{R}}_{\mathrm{alg}}[i]$ the algebraic closure of $\mathbb{Q}$}. Denote by $V_{\mathbb{C}}\subset{\mathbb{C}}^{n}$ the complexification of $V$ and choose some equations over ${\mathbb{R}}_{\mathrm{alg}}$ defining $V_{\mathbb{C}}$. Let $E|{\mathbb{Q}}$ be a finite Galois subextension of $\overline{{\mathbb{Q}}}|{\mathbb{Q}}$ containing all the coefficients of the chosen equations defining $V_{\mathbb{C}}$. By means of the action of the Galois group $\textnormal{Gal}(E|{\mathbb{Q}})$ on $V_{\mathbb{C}}$, they proved fundamental properties of real $\mathbb{Q}$}-algebraic sets and defined a precise notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity for a point of a real $\mathbb{Q}$}-algebraic set (see Definition 1.4). Latter notion of regularity over $\mathbb{Q}$} for a point on a real $\mathbb{Q}$}-algebraic set turned out to be crucial to extend the algebraic approximation techniques developed in [Nas52], [Tog73] and [AK81a]. Indeed, in [GS25] Ghiloni and the author developed such $\mathbb{Q}$}-algebraic approximation techniques and gave a complete solution of above $\mathbb{Q}$}-algebraicity problem in the case of nonsingular real algebraic sets and real algebraic sets with isolated singularities. As already mentioned, in [GS25, Theorem 1.7] Ghiloni and the author proved, as a special case, the following version over $\mathbb{Q}$} of Nash-Tognoli theorem: Every compact smooth manifold $M$ of dimension $d$ is diffeomorphic to a $\mathbb{Q}$}-nonsingular real $\mathbb{Q}$}-algebraic subset of ${\mathbb{R}}^{2d+1}$.
@@
$\mathbb{Q}$}-algebraicity problem in low dimensions: Our Main Theorem below is expected to be deeply used to provide a complete positive answer to the $\mathbb{Q}$}-algebraicity problem in low dimensions. The 1-dimensional case is just a consequence of [GS25, Theorem 1.12], in dimension $2$ and $3$ the problem is much more challenging. We refer to [BD81, AK92] for more details on the topological characterization of polyhedra of dimension up to $3$ admitting an algebraic structure and to [MP00] for necessary conditions on local invariants in all dimensions.
$\mathbb{Q}$}-algebraic homology of manifolds: It was shown by Thom [Tho54] that not every ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of compact smooth manifolds can be represented by a closed smooth submanifold, but it can be represented by the fundamental class of a compact smooth manifold by a smooth mapping. As a consequence, Kucharz [Kuc05] showed that any ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of a compact Nash manifold can be represented by a semialgebraic arc-symmetric subset. On the other hand, there are many homology classes that can not be represented by algebraic subsets (even singular), see [BD84, AK85, BK89, Ben24]. It is interesting then to see whether those classes that can be represented by real algebraic subsets can be realized by $\mathbb{Q}$}-algebraic subsets.
@@
Section 1 is devoted to review the fundamental results of $\mathbb{Q}$}-algebraic geometry developed in [FG25, GS25]. We recall the notions of $\mathbb{Q}$}-algebraic set, the decomposition in $\mathbb{Q}$}-irreducible components and the notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity of a point in a $\mathbb{Q}$}-algebraic set $V\subset{\mathbb{R}}^{n}$ (see Definition 1.2). This notion turned out to be crucial in separating the irreducible components of $\mathbb{Q}$}-algebraic sets (see Proposition 1.6) and to develop $\mathbb{Q}$}-algebraic approximation techniques in [GS25, §3]. In the last part of this first section we propose fundamental examples of $\mathbb{Q}$}-nonsingular $\mathbb{Q}$}-algebraic sets (see Definition 1.4) that became crucial in Subsection 3.1.
@@
Section 2 is devoted to the study of ${\mathbb{Z}}/2{\mathbb{Z}}$-homology groups of real Grassmannians. It is well known that each real Grassmannian ${\mathbb{G}}_{m,n}$ of affine $m$-planes in ${\mathbb{R}}^{m+n}$ can be embedded as an algebraic subset of ${\mathbb{R}}^{(m+n)^{2}}$, see [BCR98, Theorem 3.4.4]. Let us identify ${\mathbb{G}}_{m,n}$ with the above algebraic subset of ${\mathbb{R}}^{(m+n)^{2}}$. It is also well known that incidence conditions with respect to a complete flag induce a finite cellular decomposition of ${\mathbb{G}}_{m,n}$ such that the Euclidean closure of each cell is an algebraic subset of ${\mathbb{G}}_{m,n}$, see [MS74]. The closures of these cells are called Schubert varieties associated to the complete flag and they generate the ${\mathbb{Z}}/2{\mathbb{Z}}$-homology groups of each ${\mathbb{G}}_{m,n}$. The main result of this section is that, if we choose the complete flag $0\subset{\mathbb{R}}\subset{\mathbb{R}}^{2}\subset\dots\subset{\mathbb{R}}^{m+n}$ with respect to the standard coordinates of ${\mathbb{R}}^{m+n}$, each Schubert variety $\sigma_{\lambda}$ defined by incidence conditions, prescribed by a partition $\lambda$, with respect to the above complete flag admits a $\mathbb{Q}$}-desingularization (see Definition 2.3). Latter result ensures that each ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of each real Grassmannian ${\mathbb{G}}_{m,n}$ has a $\mathbb{Q}$}-nonsingular $\mathbb{Q}$}-algebraic representative. This is a key property for the construction of $\mathbb{Q}$}-algebraic relative bordisms in Section 3. We point out that very general algorithms in resolution of singularities in characteristic 0 provide a resolution which is functorial and invariant under field extension, see as references [Wł05, Section 5.7], [Kol07, Section 3.34.2] and [BM08, Theorem 1.1]. However, mentioned algorithms have a very high complexity, thus it is a very hard task in general to concretely apply them obtaining an explicit resolution of a given algebraic variety, see [FKP06]. By contrast, the application of the very explicit Bott-Samelson desingularization procedure for real embedded Schubert varieties, see [Dem74, Zel83] and the books [FP98, Man01, Bri05], allows us to conbinatorically control the equations defining the center of each blow-up of the resolution. We emphasize that, to the best of the author’s knowledge, the ${\mathbb{R}}|{\mathbb{Q}}$-regularity of real embedded Bott-Samelson varieties is not a direct consequence of results already known in the literature, although Demazure [Dem74] develops the desingularization procedure in the very general setting of flag varieties over a field $k$. Indeed, for our purposes, the choice of an appropriate embedding in some real affine space of each Grassmannian manifold ${\mathbb{G}}_{m,n}$ and of the resulting desingularization of its embedded Schubert subvarieties plays a crucial role for the notion of ${\mathbb{R}}|{\mathbb{Q}}$-regularity, see Example 1.3 for a clear explanation. We refer the interested reader to Remark 2.8 for more details on how to deduce the existence of a $\mathbb{Q}$}-desingularization of $\mathbb{Q}$}-algebraic sets from general results in resolution of singularities and why Bott-Samelson procedure is preferable in the case of Schubert varieties.
@@
Section 3 is divided in two different subsections. The first one is devoted to adapt ‘over $\mathbb{Q}$}’ the topological construction of what we call “relative bordisms” introduced by Akbulut and King in [AK81]. We stress that the topological ideas come from Akbulut and King paper but to ensure the equations to have rational coefficients, both globally and locally, our choice of an appropriate embedding is crucial. The second subsection is a review on $\mathbb{Q}$}-stable $\mathbb{Q}$}-algebraic sets introduced by Ghiloni and the author in [GS25].
@@ 1. Review on (real) $\mathbb{Q}$}-algebraic geometry
$1.$ $\mathbb{Q}$}-Algebraic sets, ${\mathbb{R}}|{\mathbb{Q}}$-regularity and $\mathbb{Q}$}-regular functions
In this subsection we briefly recall some fundamental notions of real and complex $\mathbb{Q}$}-algebraic geometry introduced in [FG25, GS25].
@@
$\mathcal{Z}_{L}(F)$ $:=\{x\in L^{n}:f(x)=0,\,\forall f\in F\},$
${\mathcal{I}}_{K}(S)$ $:=\{f\in K[x]:f(x)=0,\,\forall x\in S\}.$
@@
The $K$-dimension $\dim_{K}(V)$ of $V$ is defined as the Krull dimension of the ring $K[x]/{\mathcal{I}}_{K}(V)$. As above, $\dim_{L}(V)$ coincides with the usual dimension $\dim(V)$ of $V$, viewed as an algebraic subset of $L^{n}$. A $K$-version of a classical result concerning irreducibility and dimension holds: if $V\subset L^{n}$ is $K$-irreducible and $W\subset L^{n}$ is a $K$-algebraic set such that $W\subsetneq V$, then $\dim_{K}(W)<\dim_{K}(V)$. Moreover, if $L$ is algebraically closed or real closed, then $\dim_{K}(V)=\dim(V)$. For further information on these topics we refer to [FG25, §2].
Moreover, if $L$ is algebraically closed, an application of Hilbert’s Nullstellensatz and Galois Theory in [FG25, Theorem 2.3.4(vi)] guarantees that ${\mathcal{I}}_{L}(V)={\mathcal{I}}_{K}(V)L[x]$, thus the ideal ${\mathcal{I}}_{K}(V)$ of $K[x]$ gives complete information about $V\subset L^{n}$. Latter equality holds as well when both $L$ and $K$ are real closed fields by the Tarski-Seidenberg principle, see [FG25, Corollary 2.2.17(ii)]. However, this is false for a general field extension $L|K$. As an example, if $L|K={\mathbb{R}}|{\mathbb{Q}}$ and $V:=\{x-\sqrt[3]{2}=0\}=\{x^{3}-2=0\}\subset{\mathbb{R}}$, then
@@
We say that an algebraic set $S\subset L^{n}$ is defined over $K$ if ${\mathcal{I}}_{L}(S)={\mathcal{I}}_{K}(S)L[x]$. Clearly, if $S\subset L^{n}$ is defined over $K$, then $S\subset L^{n}$ is also $K$-algebraic. As said before, if $L$ is algebraically closed or both $L$ and $K$ are real closed, then the concepts of $K$-algebraic subset of $L^{n}$ and algebraic subset of $L^{n}$ defined over $K$ do coincide. As explained in [GS25, Appendix C], in the real algebraic setting the situation is completely different and various notions of nonsingularity over $\mathbb{Q}$} arise. In what follows we recall the definition of ${\mathbb{R}}|{\mathbb{Q}}$-regular points of a $\mathbb{Q}$}-algebraic set introduced and studied in [FG25] and whose characterizations are further investigated in [GS25].
@@
and by ${\mathcal{I}}_{\mathbb{Q}}(V)$ the vanishing ideal of $V$ in ${\mathbb{Q}}[x]$, as above. The following notion of ${\mathbb{R}}|{\mathbb{Q}}$-regular point was introduced in [FG25, Definition 5.1.1].
@@
In [FG25, Theorem 5.4.1], Fernando and Ghiloni show that the set $\operatorname{Reg}^{{\mathbb{R}}|{\mathbb{Q}}}(V)$ is a non-empty Zariski open subset of $\operatorname{Reg}(V)$. Though, it may happen that $\operatorname{Reg}^{{\mathbb{R}}|{\mathbb{Q}}}(V)\subsetneq\operatorname{Reg}(V)$, see the example below.
@@
Consider the $\mathbb{Q}$}-algebraic line $V:=\{x_{1}-\sqrt[3]{2}x_{2}=0\}=\{x_{1}^{3}-2x_{2}^{3}=0\}\subset{\mathbb{R}}^{2}$. Observe that $V$ is nonsingular (as an algebraic set), but $\overline{O}:=(0,0)$ is not ${\mathbb{R}}|{\mathbb{Q}}$-regular. Indeed, $\mathcal{I}_{{\mathbb{Q}}}(V)=(x_{1}^{3}-2x_{2}^{3})$ and $\mathcal{R}^{{\mathbb{R}}|{\mathbb{Q}}}_{V,\overline{O}}:={\mathbb{R}}[x_{1},x_{2}]_{(x_{1},x_{2})}/(x_{1}^{3}-2x_{2}^{3}){\mathbb{R}}[x_{1},x_{2}]_{(x_{1},x_{2})}$, hence $\mathcal{R}^{{\mathbb{R}}|{\mathbb{Q}}}_{V,\overline{O}}$ is not an integral domain since $x_{1}^{3}-2x_{2}^{3}=(x_{1}-\sqrt[3]{2}x_{2})(x_{1}^{2}+\sqrt[3]{2}x_{1}x_{2}+\sqrt[3]{4}x_{2}^{2})$. In particular, this implies that $\mathcal{R}^{{\mathbb{R}}|{\mathbb{Q}}}_{V,\overline{O}}$ is not a regular local ring either.
This leads to the following definition originally introduced in [GS25, Definition 1.9].